OpenStudy (anonymous):

Find the volume of a sphere with a diameter of 22 inches. Use 3.14 to approximate π. Round answers to the nearest hundredth.

ganeshie8 (ganeshie8):

use volume of sphere formula

OpenStudy (anonymous):

What is it @ganeshie8

ganeshie8 (ganeshie8):

volume = \(\large \frac{4}{3} \pi r^3\) \(r\) : radius

ganeshie8 (ganeshie8):

they gave u diameter = 22

ganeshie8 (ganeshie8):

wat wud radius be ?

OpenStudy (anonymous):

Formula?

ganeshie8 (ganeshie8):

radius will be half of diameter

ganeshie8 (ganeshie8):

radius = diameter / 2 = 22 / 2 = ?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

I'm sorry my computer has problems

ganeshie8 (ganeshie8):

its okay...

ganeshie8 (ganeshie8):

plugin \(r=11\) in the formula, and simplify

OpenStudy (anonymous):

I got it thanks